for which value of k will the pair of equations kx+ 3y= k- 3 and 12x+ ky = k have no solution? 

Grade:10

1 Answers

Arun
6 years ago

we have 

kx + 3y - k-3 = 0
12x + ky - k = 0
For given equations to have no solution
a1/a2 = b1/b2 is not equal to c1/c2
k/12 = 3/k
k^2 = 36
k = 6 or k = -6
On putting value of k in the given equation
we get a1/a2 = b1/b2 is not equal to c1/c2
hence the given equation will have no solution both for k=6 & k = -6